# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v3

@@ V-B Experimental Results
Table II presents the normalized scores of different methods. The reported scores are the normalized scores of the learned policies during the last 5 iterations, averaged over 5 random seeds. The results for all methods are obtained from the original papers. The table shows that the DOMAIN method achieves the best performance in three out of the nine datasets, comparable results in three out of the remaining six settings, and slightly poor performance in the three medium settings. Overall, DOMAIN can achieve better average performance across multiple datasets. Corresponding learning curves are shown in Fig. 5.
TABLE IV: The parameter configuration of our designed simple MDP.
@@
Figure 6: The visualization results of the DOMAIN algorithm on designed MDP. (a) illustrates the distribution of collected offline data. The intermediate color region represents the low uncertainty area, while the region on both sides represents the high uncertainty area; (b) presents the prediction results of the trained dynamic model under the initial state ${s=0}$. The red area represents the range of variance predicted by ensemble models; (c) depicts the error distribution of model data. The green color represents precise model data, while the red color represents imprecise model data; (d) illustrates the distribution of $Q$-values of model data. The black curve is the high-order fitting curve of the data. The red line indicates the optimal action, i.e. ${a\approx 0}$.
V-C Visualisation of DOMAIN Performance
@@
Due to the high dimension of states and actions in the D4RL tasks, visualizing the performance is greatly challenging. Therefore, a simple example MDP is defined for illustrative purposes following [24]. For this MDP, both the state space $\mathcal{S}$ and action space $\mathcal{A}\in[-1,1]$ are one-dimensional. The reward is defined as $R(s,a)=s$, where a larger value of state $s$ leads to a higher reward. The dynamic transition model is given by $T(s^{\prime}\mid s,a)=\mathcal{N}(\mu,\sigma)$ with the parameters shown in Table IV, indicating that the next state depends on the current action, independent of the initial state. To visualize the DOMAIN exploration in OOD region, we sample actions from $\mathcal{N}(\mu=0,\sigma=0.35)$, which means that the behavioral policy for the offline dataset is random policy.
@@
Fig. 6 (a) gives the collected offline data distribution for the above MDP, which illustrates that the actions do not cover all regions. Fig. 6 (b) presents the prediction results of the trained dynamic model. It can be concluded that the dynamic model shows higher prediction accuracy in-distribution region (low uncertain area) but lower prediction accuracy in the OOD region (high uncertain area). Fig. 6 (c) displays the model data error after 3000 iterations. It can be observed that the model data within the region exhibits relatively low errors, while the data outside the region shows larger errors. This result is consistent with the prediction results of the trained dynamic model. Fig. 6 (d) shows the model data $Q$-value curve in state ${s=0}$ after 3000 iterations. We can find that the DOMAIN algorithm does not universally underestimate all model data. Instead, it is more prone to underestimating inaccurate model data. Moreover, the value of the optimal action, i.e.,$\arg\max_{a}Q(s,a)$, is approximately equal to zero, which aligns with the selection of the offline dataset. This is because the reward is solely dependent on the state, and the state value is maximized when ${a\approx 0}$ is in-distribution region. The consistency between model data and offline data in selecting the optimal actions contributes to enhancing the stability of agent training.
@@
TABLE V: The analysis results for parameter $H$ and $\lambda$. Bold indicates the highest score in ablation experiments. Highlighted numbers indicate the optimal results in Table III.
@@ V-D The Parameter Analysis
This part gives further analysis for weight $\lambda$ and rollout length $H$. Table V analyzes the impact of two parameter changes on the performance of the agent. Due to high training cost, we fix one parameter and test the impact of changes in another parameter on agent performance.
@@
The effect of weight $\lambda$. The parameter $\lambda$ is used to regularization between Bellman error and $Q$ value of data, and guarantees the lower bound of $Q$-function for inaccuracy model data. A larger $\lambda$ is needed to guarantee the lower bound of inaccuracy model data in narrower data distribution. However, for diverse data distribution, too large $\lambda$ may cause the lower bound to be too loose, thus reducing agent performance. Table V shows that in narrower “Medium” and “Medium-Expert” datasets, too small $\lambda$ leads to unstable training, thus reducing the agent performance. For diverse “Medium-Replay” datasets, too large $\lambda$ causes the high constraint, thus limiting the improvement of the agent.
@@
The effect of rollout length $H$. When rollout length $H=1$, the trained environment model has high prediction accuracy. As $H$ increases, the accuracy of prediction decreases, but it is beneficial to explore the OOD region. If the model data error is too large, the algorithm will not be able to achieve a balance between accurate offline data and inaccurate model data, leading to unstable agent training. Therefore, it is necessary to ensure that the model data error is not too large. Table V shows that, as $H$ increases, the agent performance shows a downward trend under the Walker2d task since the error of the trained environment is large (Fig. 3). For Halfcheetah and Hopper tasks, the environment error is small. Therefore, as $H$ increases, the agent performance does not show a downward trend, but the optimal $H$ is different for different datasets.
VI Conclusion
@@
This paper proposes a mildly conservative model-based offline RL (DOMAIN) algorithm. By designing a model data adaptive sampling distribution, DOMAIN can dynamically adjust the model data penalty, increasing the penalty for inaccuracy model data while reducing the penalty for accurate model data, thereby reducing its conservatism. Moreover, DOMAIN avoids unreliable estimation of model uncertainty. Theoretical analysis demonstrates that DOMAIN can guarantee safe policy improvements and $Q$-values learned by DOMAIN outside the region are lower bounds of the true $Q$-values. Experimental results on the D4RL benchmark show that DOMAIN outperforms previous RL algorithms on most datasets. Since incorporating conservatism during the training process constrains the improvement of agent performance, future efforts can focus on optimizing the environment model itself to further enhance the overall performance.
Appendix A Proof of Relevant Theorems
A-A Proof of Theorem 1
@@
$\widehat{\mathcal{T}}^{\pi}Q(s,a)$ $=(1-f)\left\{r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q\left(s^{\prime},a^{\prime}\right)\right\}$ (18)
$+f\left\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]\right\}$
@@
$\|\widehat{\mathcal{T}}^{\pi}Q_{1}-\widehat{\mathcal{T}}^{\pi}Q_{2}\|_{\infty}$ (19)
$=$ $(1-f)\gamma\max_{s,a}\left|\max_{a^{\prime}\in\mathcal{A}}Q_{1}\left(s^{\prime},a^{\prime}\right)-\max_{a^{\prime}\in\mathcal{A}}Q_{2}\left(s^{\prime},a^{\prime}\right)\right|$
$+f\gamma\max_{s,a}\left|\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q_{1}]-\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q_{2}]\right|$
$\leq$ $(1-f)\gamma\max_{s,a}\left|\max_{a^{\prime}\in\mathcal{A}}Q_{1}\left(s^{\prime},a^{\prime}\right)-\max_{a^{\prime}\in\mathcal{A}}Q_{2}\left(s^{\prime},a^{\prime}\right)\right|$
$+f\gamma\max_{s,a}\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}\left|[\max_{a^{\prime}\in\mathcal{A}}Q_{1}-\max_{a^{\prime}\in\mathcal{A}}Q_{2}]\right|$
@@
According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$. Thus, Theorem 1 is proven.
A-B Proof of Theorem 2
@@
$\left|r_{\mathcal{M}}(s,a)-r_{\overline{\mathcal{M}}}(s,a)\right|<\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}$ (20)
$\left|T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)-T_{\overline{\mathcal{M}}}\left(s^{\prime}\mid s,a\right)\right|_{1}<\frac{C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}$
@@
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|<\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}}$
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|<\varepsilon_{r}+D\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\frac{2\gamma R_{\max}}{1-\gamma}$
Proof. The difference between the Bellman operator $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ under the empirical MDP and the actual Bellman operator $\mathcal{T}_{\mathcal{M}}^{\pi}$ is given by:
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|$ (21)
$\leq$ $\Big{|}r_{\mathcal{M}}(s,a)-r_{\overline{\mathcal{M}}}(s,a)\Big{|}+\gamma\Big{|}\sum_{s^{\prime}}(T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right).$
$-T_{\overline{\mathcal{M}}}\left(s^{\prime}\mid s,a\right))\mathbb{E}_{a^{\prime}\sim\pi}\left[\widehat{Q}^{k}\left(s^{\prime},a^{\prime}\right)\right]\Big{|}$
@@
$\left|\left(\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}\right)-\left(\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right)\right|$ (22)
$\leq$ $\Big{|}r_{\mathcal{M}}(s,a)-r_{\widehat{\mathcal{M}}}(s,a)\Big{|}+\gamma\Big{|}\sum_{s^{\prime}}(T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)$
$-T_{\widehat{\mathcal{M}}}\left(s^{\prime}\mid s,a\right))\mathbb{E}_{a^{\prime}\sim\pi}\left[\widehat{Q}^{k}\left(s^{\prime},a^{\prime}\right)\right]\Big{|}$
@@
The proof derivation of Theorem 2 is presented below. First, we introduce the following notation: Let $\mathcal{T}^{\pi}Q=r+\gamma P^{\pi}Q$, where $P^{\pi}Q(s,a)=\mathbb{E}_{s^{\prime}\sim T(\cdot\mid s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q^{\pi}(s^{\prime},a^{\prime})]$. Based on Eq. (14) and the Lemma 1, it can obtain:
$\widehat{Q}^{k+1}=(1-f)\mathcal{T}_{\overline{\mathcal{M}}}^{{\pi}}\widehat{Q}^{k}+f\widehat{\mathcal{T}}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}-\lambda\eta(s,a)$ (23)
$\leq$ $\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}+(1-f)\underbrace{\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}}}_{:=\Delta_{1}(s,a)}-\lambda\eta(s,a)$
$+f(\underbrace{\left|r_{\mathcal{M}}-r_{\widehat{\mathcal{M}}}\right|+D\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\frac{2\gamma R_{\max}}{1-\gamma}}_{:=\Delta_{2}(s,a)})$
@@
By applying the fixed-point theorem, the above equation leads to:
$\widehat{Q}^{\pi}$ $\leq\mathcal{T}^{\pi}\widehat{Q}^{\pi}+\Delta=r+\gamma P^{\pi}\widehat{Q}^{\pi}+\Delta$ (24)
$\Rightarrow\widehat{Q}^{\pi}$ $\leq\left(I-\gamma P^{\pi}\right)^{-1}(r+\Delta)$
$=\left(I-\gamma P^{\pi}\right)^{-1}r+\left(I-\gamma P^{\pi}\right)^{-1}\Delta$
@@
$\widehat{Q}^{\pi}(s,a)\leq Q^{\pi}(s,a)+\left(I-\gamma P^{\pi}\right)^{-1}\Delta(s,a)$ (25)
@@
$\widehat{V}^{\pi}(s)\leq V^{\pi}(s)+\left(I-\gamma P^{\pi}\right)^{-1}\mathbb{E}_{a\sim\pi(a\mid s)}[\Delta(s,a)]$ (26)

### v4

@@ VI-B Experimental Results
Table III presents the normalized scores of different methods. The reported scores are the normalized scores of the learned policies during the last 5 iterations, averaged over 5 random seeds. The results for all methods are obtained from the original papers. Due to the differences between datasets under different tasks, RL methods hardly perform well on all datasets. This table shows that our method DOMAIN performs poorly on the “Medium” dataset, ARMOR [37] performs poorly on the “Medium-replay” dataset, and RAMBO [36] performs poorly on the “Medium-expert” dataset. This table also indicates that the DOMAIN method achieves the best performance in three out of the nine datasets, and comparable results in three out of the remaining six settings. and a slightly poor performance in the three medium settings. Overall, DOMAIN can achieve better average performance across multiple datasets. Corresponding learning curves are shown in Fig. 5.
The reason why DOMAIN performs poorly in “Medium” dataset is that the distribution of medium is narrower than that of medium-replay and medium-expert. Then, the model data generated by the environment model trained based on the dataset is also narrow, and the trained discriminator hardly distinguishes model data from the offline dataset. As a result, the adaptive sampling distribution $\omega(s,a)$ is inaccurate, which reduces the final agent performance. Moreover, Table III shows that the performance of the “Medium-expert” dataset in three tasks is almost always superior to that of “Medium” and “Medium-replay” datasets for any algorithms. The difference between “Medium” and “Medium-replay” is that “Medium-replay” contains more data at different levels, which can facilitate agent exploration. However, the algorithm may not necessarily perform better than “Medium” under “Medium-replay” for different tasks. This mainly depends on the adaptability of the algorithm itself to diverse data and the coverage of the data.
TABLE V: The parameter configuration of our designed simple MDP.
@@
TABLE VI: Comparison results of computation time (unit: seconds).
Method Environment Average
Walker2d Hopper Halfcheetah
COMBO $44422.1$ $41477.9$ $49271.5$ $43908.6$
RAMBO $91853.9$ $86531.3$ $96766.3$ $91717.2$
DOMAIN $46957.7$ $46930.9$ $48244.5$ $47377.7$
The computational complexity is an important index for the performance of the proposed method. The training time of model-based offline RL usually includes two parts: environment learning and policy learning. Table VI compares computation cost between DOMAIN and the current two typical model-based offline RL methods (COMBO [28] and RAMBO [36]) under the three environments. This table shows that DOMAIN can reduce computational costs by nearly half compared to RAMBO, primarily because the environment model is continuously updated during the policy learning process in RAMBO. The computational complexity of DOMAIN is slightly higher than that of COMBO, as DOMAIN requires the training of a discriminator to solve the adaptive sampling distribution $\omega$, which introduces additional training costs. In general, compared with the above two methods, DOMAIN can achieve better control performance with relatively small computational complexity.
@@
Figure 6: The visualization results of the DOMAIN algorithm on the designed MDP. (a) illustrates the distribution of collected offline data. The intermediate color region represents the low uncertainty area, while the region on both sides represents the high uncertainty area; (b) presents the prediction results of the trained dynamic model under the initial state ${s=0}$. The red area represents the range of variance predicted by ensemble models; (c) depicts the error distribution of model data. The green color represents precise model data, while the red color represents imprecise model data; (d) illustrates the distribution of $Q$-values of model data. The black curve is the high-order fitting curve of the data. The red line indicates the optimal action, i.e. ${a\approx 0}$.
VI-C Visualization of DOMAIN Performance
@@
Due to the high dimension of states and actions in the D4RL tasks, visualizing the performance is greatly challenging. Therefore, a simple example MDP is defined for illustrative purposes following [36]. For this MDP, both the state space $\mathcal{S}$ and action space $\mathcal{A}\in[-1,1]$ are one-dimensional. The reward is defined as $R(s,a)=s$, where a larger value of state $s$ leads to a higher reward. The dynamic transition model is given by $T(s^{\prime}\mid s,a)=\mathcal{N}(\mu,\sigma)$ with the parameters shown in Table V, indicating that the next state depends on the current action, independent of the initial state. To visualize the DOMAIN exploration in the OOD region, we sample actions from $\mathcal{N}(\mu=0,\sigma=0.35)$, which means that the behavioral policy for the offline dataset is a random policy.
@@
Fig. 6 (a) shows the distribution of the collected offline data for the aforementioned MDP, illustrating that the actions do not cover all regions. Fig. 6 (b) presents the prediction results of the trained dynamic model, which demonstrate higher accuracy in the in-distribution region (low-uncertainty area) and lower accuracy in the OOD region (high-uncertainty area). Fig. 6 (c) displays the model data error after 3000 iterations. It can be observed that the model data within the region exhibits relatively low errors, while the data outside the region shows larger errors. This result is consistent with the prediction results of the trained dynamic model. Fig. 6 (d) shows the model data $Q$-value curve in state ${s=0}$ after 3000 iterations. We can find that the DOMAIN algorithm does not universally underestimate all model data. Instead, it is more prone to underestimating inaccurate model data. Moreover, the value of the optimal action, i.e.,$\arg\max_{a}Q(s,a)$, is approximately zero, which aligns with the selection of the offline dataset. This is because the reward is solely dependent on the state, and the state value is maximized when ${a\approx 0}$ is in-distribution region. The consistency between model data and offline data in selecting the optimal actions contributes to enhancing the stability of agent training.
VI-D Robust Experiments
The above experiments are based on a simulation environment. However, there is a gap between the simulation and the actual environment, and there are different types of disturbances in the practical scenario. Therefore, it is a necessity to test the performance of the proposed method in the above situations. Here, we conduct two types of experiments: one is the simulation environment parameter perturbation experiment, and another is the external disturbance experiment.
Environment parameter perturbations: Achieving high-fidelity simulation environments is often challenging, which leads to inherent sim-to-real gaps in policy learning. In the case of D4RL tasks, the simulation environment relies on the MuJoCo physics engine. To assess the performance of our method in environments with large dynamics gaps, we modify key MuJoCo parameters, including friction and gravity, to create an environment with a substantial sim-to-real gap. Fig. 7 (a) and (b) present the results under various test conditions, where different gravity and friction coefficients are applied to the environment, with darker colors indicating superior performance. This figure also illustrates that under halfcheetah, variations in gravity have a significant impact on performance. Conversely, under hopper, changes in friction exert a substantial influence on performance.
Figure 7: The results of robust experiments for DOMAIN. (a) and (b) give the performance under environment parameter random perturbations in halfcheetah and hopper environments, respectively.The gravity and friction vary from 0.5 to 5 times and 0.5 to 1.5 times the value of the simulation environment, respectively. (c) and (d) show the performance under attack scales range $[0,0.2]$ of RA and AD two attack types in walker2d, hopper and halfcheetah environments, respectively
External disturbances: Due to the observation errors and external disturbances in real-world scenarios, the robustness of the policy is evaluated using the two perturbation methods below, referred to as RA and AD [14, 15]. For RA method, the perturbed states are randomly sampled from the set $U_{\epsilon}(s)=\{s^{\prime}\in\mathcal{S}:d(s,s^{\prime})^{p}\leq\epsilon\}$, where $\epsilon$ denotes the scale of perturbation and $d(s,s^{\prime})^{p}$ measures the distance between $s$ and $s^{\prime}$ [14]. AD is an effective perturbation method based on the agent’s policy, which selects the state $s^{\prime}$ from $U_{\epsilon}(s)$ that maximizes the gap between action distribution $\pi_{\theta}(\cdot|s)$ and $\pi_{\theta}(\cdot|s^{\prime})$, i.e. $\min_{s^{\prime}\in U_{\epsilon}(s)}-D_{J}(\pi_{\theta}(\cdot|s)||\pi_{\theta}(\cdot|s^{\prime}))$ [40]. The observation states are attacked through the above perturbation in the evaluation phase. Fig. 7 (c) and (d) give the result of DOMAIN with varying perturbation scale $\epsilon$ under RA and AD attack in three environments. As external disturbances progressively increase, the performance exhibits a declining trend.
Based on the experimental results presented in Fig. 7, it is evident that DOMAIN is capable of adapting to minor external disturbances and variations in environmental parameters. However, its performance significantly deteriorates under large-scale variations. Future research should focus on enhancing the robustness of the algorithm to address these limitations. For example, the robust Bellman operator can be introduced into the standard Bellman operator, or the idea of the robust adversarial model can be introduced into model training.
@@ VI-D Robust Experiments
TABLE VII: The agent performance results averaged over 5 seeds under different parameters $H$ and $\lambda$. Bold indicates the highest score in ablation experiments. Highlighted numbers indicate the optimal results in Table IV.
@@ VI-E The Parameter Analysis
This part gives further analysis for weight $\lambda$ and rollout length $H$. Note that the tuning results in Table V are not optimal with high probability. Table VII analyzes the impact of two parameter changes on the performance of the agent. Due to high training cost, we fix one parameter and test the impact of changes in another parameter on agent performance. Note that the tuning results presented in Table IV are not optimal with high probability. The optimal parameters can be found through more precise and wide-ranging parameter searches, but this requires significant training costs. The parameter ranges specified in Table VII are primarily intended to analyze the effect of parameter variations on agent performance.
@@
The effect of weight $\lambda$. The parameter $\lambda$ is used to regularization between Bellman error and $Q$ value of data, and guarantees the lower bound of $Q$-function for inaccuracy model data. A larger $\lambda$ is needed to guarantee the lower bound of inaccuracy model data in a narrower data distribution. However, for diverse data distribution, too large $\lambda$ may cause the lower bound to be too loose, thus reducing agent performance. Table VII shows that in narrower “Medium” and “Medium-Expert” datasets, too small $\lambda$ leads to unstable training, thus reducing the agent’s performance. For diverse “Medium-Replay” datasets, too large $\lambda$ causes the high constraint, thus limiting the improvement of the agent.
@@
The effect of rollout length $H$. When rollout length $H=1$, the trained environment model has high prediction accuracy. As $H$ increases, the accuracy of prediction decreases, but it is beneficial to explore the OOD region. If the model data error is too large, the algorithm will not be able to achieve a balance between accurate offline data and inaccurate model data, leading to unstable agent training. Therefore, it is necessary to ensure that the model data error is not too large. Table VII shows that, as $H$ increases, the agent’s performance shows a downward trend under the Walker2d task since the error of the trained environment is large (Fig. 3). For Halfcheetah and Hopper tasks, the environment error is small. Therefore, as $H$ increases, the agent’s performance does not show a downward trend, but the optimal $H$ is different for different datasets.
VII Conclusion
@@
This paper proposed a mildly conservative model-based offline RL (DOMAIN) algorithm. By designing a model data adaptive sampling distribution, DOMAIN can dynamically adjust the model data penalty, increasing the penalty for inaccuracy model data while reducing the penalty for accurate model data, thereby reducing its conservatism. Moreover, DOMAIN avoided unreliable estimation of model uncertainty. Theoretical analysis demonstrated that DOMAIN can guarantee safe policy improvements and $Q$-values learned by DOMAIN outside the region are lower bounds of the true $Q$-values. Experimental results on the D4RL benchmark showed that DOMAIN outperforms previous RL algorithms on most datasets. Despite the promising results, this study presents several limitations that warrant further investigation:
1) The incorporation of conservatism during the training process inherently restricts the potential performance improvement of the agent. Future work could explore the optimization of the environment model itself to achieve enhanced agent performance without compromising safety constraints.
2) The quality of model-generated data is highly dependent on the rollout length, which was empirically determined for different tasks through manual tuning in this study. To improve the generalizability of the proposed method, subsequent research should investigate an adaptive rollout strategy that dynamically adjusts the rollout length based on real-time data quality metrics.
3) The experimental validation of the proposed algorithm has been confined to simulation environments. To further demonstrate its practical applicability, future research should include offline data collection from physical robotic locomotion tasks, followed by comprehensive network training and policy deployment in real-world scenarios.
Supplementary Materials (Appendix)
This section provides the detailed proof for the convergence of Bellman operator, the lower bound for value function, mildly conservative estimation of $Q$-value and policy improvement guarantee (Theorem 1- 4).
VII-A Proof of Theorem 1
@@
$\widehat{\mathcal{T}}^{\pi}Q(s,a)$ $=(1-f)\left\{r(s,a)+\gamma\max_{a^{\prime}\in\mathcal{A}}Q\left(s^{\prime},a^{\prime}\right)\right\}+f\left\{r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]\right\}$ (S.1)
@@
$\|\widehat{\mathcal{T}}^{\pi}Q_{1}-\widehat{\mathcal{T}}^{\pi}Q_{2}\|_{\infty}$ (S.2)
$=$ $(1-f)\gamma\max_{s,a}\left|\max_{a^{\prime}\in\mathcal{A}}Q_{1}\left(s^{\prime},a^{\prime}\right)-\max_{a^{\prime}\in\mathcal{A}}Q_{2}\left(s^{\prime},a^{\prime}\right)\right|+f\gamma\max_{s,a}\left|\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q_{1}]-\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}[\max_{a^{\prime}\in\mathcal{A}}Q_{2}]\right|$
$\leq$ $(1-f)\gamma\max_{s,a}\left|\max_{a^{\prime}\in\mathcal{A}}Q_{1}\left(s^{\prime},a^{\prime}\right)-\max_{a^{\prime}\in\mathcal{A}}Q_{2}\left(s^{\prime},a^{\prime}\right)\right|+f\gamma\max_{s,a}\mathbb{E}_{s^{\prime}\sim T_{\widehat{\mathcal{M}}}}\left|[\max_{a^{\prime}\in\mathcal{A}}Q_{1}-\max_{a^{\prime}\in\mathcal{A}}Q_{2}]\right|$
@@
According to the fixed point theorem, any initial $Q$ function can converge to a unique fixed point $Q^{*}$ by repeatedly applying the Bellman operator $\widehat{\mathcal{T}}^{\pi}$. This completes the proof of Theorem 1.
VII-B Proof of Theorem 2
@@
$\left|r_{\mathcal{M}}(s,a)-r_{\overline{\mathcal{M}}}(s,a)\right|<\frac{C_{r,\delta}}{\sqrt{|\mathcal{D}(s,a)|}},\qquad\left|T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)-T_{\overline{\mathcal{M}}}\left(s^{\prime}\mid s,a\right)\right|_{1}<\frac{C_{T,\delta}}{\sqrt{|\mathcal{D}(s,a)|}}$ (S.3)
@@
$\left|\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}-\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right|<\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}},\qquad\left|\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}-\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right|<\varepsilon_{r}+D\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\frac{2\gamma R_{\max}}{1-\gamma}$
Proof for Lemma 1. The difference between the Bellman operator $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ under the empirical MDP and the actual Bellman operator $\mathcal{T}_{\mathcal{M}}^{\pi}$ is given by:
$\left|\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}-\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right|\leq$ $\Big{|}r_{\mathcal{M}}(s,a)-r_{\overline{\mathcal{M}}}(s,a)\Big{|}+\gamma\Big{|}\sum_{s^{\prime}}(T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)-T_{\overline{\mathcal{M}}}\left(s^{\prime}\mid s,a\right))\mathbb{E}_{a^{\prime}\sim\pi}\left[\widehat{Q}^{k}\left(s^{\prime},a^{\prime}\right)\right]\Big{|}$ (S.4)
@@
$\left|\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}-\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}\right|\leq$ $\Big{|}r_{\mathcal{M}}(s,a)-r_{\widehat{\mathcal{M}}}(s,a)\Big{|}+\gamma\Big{|}\sum_{s^{\prime}}(T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)-T_{\widehat{\mathcal{M}}}\left(s^{\prime}\mid s,a\right))\mathbb{E}_{a^{\prime}\sim\pi}\left[\widehat{Q}^{k}\left(s^{\prime},a^{\prime}\right)\right]\Big{|}$ (S.5)
@@
Proof for Theorem 2. First, we introduce the following notation: Let $\mathcal{T}^{\pi}Q=r+\gamma P^{\pi}Q$, where $P^{\pi}Q(s,a)=\mathbb{E}_{s^{\prime}\sim T(\cdot\mid s,a),a^{\prime}\sim\pi(\cdot\mid s^{\prime})}[Q^{\pi}(s^{\prime},a^{\prime})]$. Based on Eq. (14) and the Lemma 1, it can obtain:
$\widehat{Q}^{k+1}=(1-f)\mathcal{T}_{\overline{\mathcal{M}}}^{{\pi}}\widehat{Q}^{k}+f\widehat{\mathcal{T}}_{\widehat{\mathcal{M}}}^{\pi}\widehat{Q}^{k}-\lambda\eta(s,a)$ (S.6)
$\leq$ $\mathcal{T}_{\mathcal{M}}^{\pi}\widehat{Q}^{k}+(1-f)\underbrace{\frac{C_{r,T,\delta}R_{\max}}{(1-\gamma)\sqrt{|\mathcal{D}(s,a)|}}}_{:=\Delta_{1}(s,a)}-\lambda\eta(s,a)+f(\underbrace{\left|r_{\mathcal{M}}-r_{\widehat{\mathcal{M}}}\right|+D\left(T_{\mathcal{M}},T_{\widehat{\mathcal{M}}}\right)\frac{2\gamma R_{\max}}{1-\gamma}}_{:=\Delta_{2}(s,a)})$
@@
By applying the fixed-point theorem, the above equation yields the following result:
$\widehat{Q}^{\pi}$ $\leq\mathcal{T}^{\pi}\widehat{Q}^{\pi}+\Delta=r+\gamma P^{\pi}\widehat{Q}^{\pi}+\Delta$ (S.7)
$\Rightarrow\widehat{Q}^{\pi}$ $\leq\left(I-\gamma P^{\pi}\right)^{-1}(r+\Delta)=\left(I-\gamma P^{\pi}\right)^{-1}r+\left(I-\gamma P^{\pi}\right)^{-1}\Delta$
@@
$\widehat{Q}^{\pi}(s,a)\leq Q^{\pi}(s,a)+\left(I-\gamma P^{\pi}\right)^{-1}\Delta(s,a)$ (S.8)
@@
$\widehat{V}^{\pi}(s)\leq V^{\pi}(s)+\left(I-\gamma P^{\pi}\right)^{-1}\mathbb{E}_{a\sim\pi(a\mid s)}[\Delta(s,a)]$ (S.9)